### v1

# A deep reinforcement learning agent trained for interval timing exhibits similarities to biological systems

Abstract
Drawing parallels between Deep Artificial Neural Networks (DNNs) and biological systems can aid in understanding complex biological mechanisms that are difficult to disentangle. Temporal processing, an extensively researched topic, is one such example that lacks a coherent understanding of its underlying mechanisms. In this study, we investigate temporal processing in a Deep Reinforcement Learning (DRL) agent performing an interval timing task and explore potential biological counterparts to its emergent behavior. The agent was successfully trained to perform a duration production task, which involved marking successive occurrences of a target interval while viewing a video sequence. Analysis of the agent’s internal states revealed oscillatory neural activations, a ubiquitous pattern in biological systems. Interestingly, the agent’s actions were predominantly influenced by neurons exhibiting these oscillations with high amplitudes. Parallels are drawn between the agent’s time-keeping strategy and the Striatal Beat Frequency (SBF) model, a biologically plausible model of interval timing. Furthermore, the agent maintained its oscillatory representations and task performance when tested on different video sequences (including a blank video). Thus, once learned, the agent internalized its time-keeping mechanism and showed minimal reliance on its environment to perform the timing task. A hypothesis about the resemblance between this emergent behavior and certain aspects of the evolution of biological processes like circadian rhythms, has been discussed. This study aims to contribute to recent research efforts of utilizing DNNs to understand biological systems, with a particular emphasis on temporal processing.
Introduction
A recent research interest has emerged in drawing parallels between Deep Artificial Neural Networks (DNNs) and biological systems (Kanwisher et al.,, 2023; Richards et al.,, 2019). The similarities between the two are noteworthy because, DNNs exhibit emergent behavior and are not explicitly engineered to mimic biological systems. Unlike biological systems, which are a result of complex evolutionary adaptations, DNNs rely on relatively simple optimization techniques and architectures to solve problems. These two distinct systems converging to similar behaviors suggests important implications about the nature of the problem being solved. It may indicate that certain behaviors and representations are fundamentally useful to find a solution, and that problems presumed to require complex solutions might, in fact, be addressed through simpler means. Additionally, DNNs, with their flexibility for ablation studies and architectural manipulations, provide a powerful platform for systematically testing different hypotheses. Hence, identifying similarities between the two systems can facilitate better understanding of biological systems.
Temporal processing is an extensively researched topic in both psychology and neuroscience (Wittmann,, 2009). However, the lack of a dedicated sensory organ for perceiving time, and scattered evidence regarding the involvement of multiple brain regions, makes it difficult to form a coherent understanding of temporal processing in humans and animals. Interval timing provides a framework to study these temporal processing mechanisms. The ability of humans and animals to successfully produce, reproduce, or categorize different temporal durations suggests the existence of an internal time-keeping mechanism. Recent research has begun exploring interval timing also in deep reinforcement learning (DRL) agents (Deverett et al.,, 2019; Lin et al.,, 2023). These studies additionally explore potential biological counterparts to the internal time-keeping mechanisms that emerge in the DRL agent performing interval timing tasks. They put forth a discussion about using advancements in Artificial Intelligence (AI) to enhance our understanding of temporal encodings in the brain.
In this study, we investigate the internal time-keeping mechanisms of a DRL agent trained on a duration production task (Thönes and Stocker,, 2019). We show that the agent develops a robust internal time-keeping mechanism, characterized by oscillatory activation patterns, a motif commonly observed in biological systems. Furthermore, the agent’s timing strategy shares key features with the Striatal Beat Frequency (SBF) model, a biologically plausible model of interval timing (Matell and Meck,, 2000). Finally, the robust and internalized time-keeping mechanism developed in the agent is compared to certain aspects of the evolution of biological processes like circadian rhythms. Through this study, we aim to contribute to research that draws parallels between temporal processing in artificial agents and biological systems. It is our view that such efforts may provide alternative approaches to understanding temporal processing in humans and animals.
Methodology
The interval timing task used in this study is a time production task that involved watching a video sequence and successively marking a fixed time interval. The DRL agent received one frame at a time from its environment. Thus, the agent’s state (observation) at a given point was the current frame it was viewing. These frames were a part of a video sequence that was a randomly chosen sequence of twenty frames extracted from a single video. Thus, one episode comprised of twenty frames. The agent had two possible actions: ”Go”, to proceed to the next video frame and ”Interval”, to mark an interval (i.e., produce the target interval) and proceed to the next frame. When the agent produced the correct interval, it received a reward of +1; for incorrect productions or missed intervals, it received a penalty of -1; all other cases generated a reward of 0. Different agents were trained to produce target intervals of two, three, four and five time steps. In the specific examples used in this study, the agent was trained to produce four time steps. Thus, the agent had to select the action ”Interval” on every fourth frame. The important findings discussed in this study are consistent across agents trained for other target intervals.
The DRL agent architecture consisted of three primary components: a Convolutional Neural Network (CNN) to extract spatial features from each video frame, a Long Short-Term Memory (LSTM) network to process temporal dependencies across frames, and a linear layer to process the resulting representation and to compute the state value and next action (see Figure 1). The agent was implemented using the recurrent policy variant of the Proximal Policy Optimization (PPO) algorithm provided by the Stable-Baselines3 (SB3) library (Raffin et al.,, 2021). The reinforcement learning environment was a custom built environment developed using the Gymnasium framework (Towers et al.,, 2024). It was designed to deliver sequential video frames and handle action and rewards specific to the target interval time.
The agent was evaluated on a set of video sequences extracted from videos other than the training video. These test videos were categorized based on their Mean Absolute Frame Difference (MAFD), computed as described in Salih and George, (2020) and averaged over all frames for each video sequence (see Table 1). We use the average MAFD as a simplified metric to quantify the amount of information in a video sequence. A blank video, or one in which the same frame is repeated throughout (i.e., constant frame video), yields a MAFD of zero and thus contains minimal information. In contrast, videos with changing frames exhibit higher MAFD values and thus contain more information.
Figure 1: Overview of the DRL agent architecture and input processing pipeline: The DRL agent’s architecture consists of an action network (left) and a value network (right), both processing the input frames in parallel. The agent receives one frame at each time step from the environment. The input shape transitions, from the original frame to the action and state value, are shown in parentheses. At each time step, the agent can take one of the two actions: “GO” or “Interval,” to advance to the next frame and receive a reward from the environment.
Neural oscillations - an emergent time-keeping mechanism in the artificial agent
The agent successfully managed to learn the timing task and was able to produce the target interval consecutively throughout the video sequence. Analysis of the agent’s internal states revealed that the LSTM layer played a central role in the development of an intrinsic timekeeping mechanism. This was expected since the LSTM network, which is a type of artificial Recurrent Neural Network (RNN), is designed to extract temporal features from sequential data (Hochreiter and Schmidhuber,, 1997; Yu et al.,, 2019). The hidden state activations of the LSTM layer were passed to the action network which was responsible for selecting the next action. Thus, the agent primarily relied on its LSTM layer to determine its next action and successfully perform the timing task.
A Fourier Transform of LSTM hidden state activations revealed a clear oscillatory pattern. Figure 2 shows the Fast Fourier Transform (FFT) of the hidden state activations over time. It shows that the majority of neurons exhibit oscillations at a frequency of 0.25, corresponding to one peak every four time steps. Notably, neurons with high weight magnitudes (z-score >> 2) for the “Interval” action in the action network (highlighted in blue in Figure 2) show greater oscillation amplitudes compared to those with lower weight magnitudes (z-score << -2) for both actions (highlighted in red in Figure 2). This suggests that the agent’s action decisions are more strongly influenced by neurons exhibiting higher oscillatory amplitudes.
A Principal Component Analysis (PCA) was conducted to identify dominant patterns in the activations of the 256 hidden state neurons in the LSTM network across time steps. The first two principal components that explained 86% of variance, further verified the occurrence of oscillatory patterns across time in the hidden state neural activations (see Figure 3). Interestingly, a prior study involving a DRL agent trained for time reproduction (with spatial input) did not report such oscillatory patterns in the PCA of hidden states (Deverett et al.,, 2019). The authors concluded that the hidden units in the RNN layer of the agent exhibit counter-like behavior. However, in Lin et al.,, 2023, for certain conditions, oscillating neurons have been reported in the action and value network of a DRL agent (with non-spatial inputs) for a duration comparison and an episodic working memory task. This suggests that oscillatory dynamics may serve as a fundamental mechanism of time-keeping for specific types of interval timing tasks in artificial agents.
Emergence of oscillatory patterns in the DRL agent’s network is noteworthy because oscillations are a known characteristic of various types of biological neurons (Heltberg et al.,, 2021; Buzsáki et al.,, 2013), including those involved in temporal processing. Importantly, the DRL network was not provided with any external timing cues or explicitly encoded neural dynamics. The agent’s only inputs were reward signals and the next video frame from its environment. Despite this, it successfully learned to perform the timing task, and, remarkably, the hidden state activations of the network converged to an oscillatory pattern. This suggests that such dynamics may emerge naturally as a functional solution for time-keeping in specific interval timing task.
Figure 2: Fast Fourier Transform (FFT) of the hidden state activations over time. The blue lines represent neurons with high weight magnitudes (z-score >> 2) for the “Interval” action in the action network , while the red lines represents neurons with low weight magnitudes (z-score << -2) for both actions. The gray lines correspond to all other neurons. All neurons exhibit oscillation with a frequency of 0.25 (1 peak every 4 time step).
Figure 3: Principal component analysis (PCA) of the hidden state neural activations. The first and second components (in dark and light blue respectively) which exhibit an oscillatory pattern across time steps, explain 86% of variance.
Parallels with biologically plausible timing models
A biologically plausible model of interval timing called the Striatal Beat Frequency model (SBF), describes interval timing in terms of coincidence detection in neural oscillations Matell and Meck, (2000). According to this model, the striatal spiny neurons function as coincidence detectors of cortical neural oscillations. These coincidence detectors fire when a set amount of coincidental input activity (i.e., coincident oscillations of the cortical neurons) is reached. According to Buhusi and Meck, (2005) and Matell and Meck, (2000), for tasks involving the production of a learned interval, there is a burst of dopaminergic activity at the trial onset. This triggers synchronization in oscillation of the cortical neurons. Additionally, another burst in activity of dopaminergic neurons occurs at the expected time of reward which corresponds to the target interval. This causes an update in the weights of corticostriatal synapses and consequently triggers the firing of the striatal spiny neurons (or the coincidental detectors). As a result, the firing rate of these coincidence detectors can encode information about a previously learned interval.
Figure 4 shows the raw activations of the 256 hidden state neurons across time steps. Neurons highlighted in blue represent high weight magnitudes (z-score >> 2) for the “Interval” action in the action network, indicating a strong influence on the decision of the next action. On the other hand, neurons highlighted in red correspond to low weight magnitudes (z-score << -2) for both actions in the action network, implying a weaker influence on the action selection. The activations of the blue neurons show a high amplitude oscillatory pattern across time steps with a peak at the target interval (i.e., every fourth time step). Thus, the neural activations of these neurons peak at the time steps corresponding to reward delivery (all other time steps are either related to no reward - for a correct action, or penalty - for an incorrect action). On the other hand, the activations of red neurons show a trough at the the target interval.
The emergence of neural activations that peak at reward delivering time steps and the high weights assigned by the action network to neurons with a high positive peak, bear resemblance to the functioning of the SBF model. In this analogy, the action network acts like a coincidence detector, producing an action of marking an interval when the activations of the (oscillating) neurons cross a learned threshold.
Figure 4: Activations of the 256 hidden state neurons in the LSTM layer of the DRL agent over time. Neurons shown in blue had high weight magnitudes (z-score >> 2.0) for the “Interval” action in the agent’s action network. While those in red had the least magnitudes (z-score << -2.0) for both actions. Thus, neurons with high amplitude oscillating patterns had the highest contribution in action selection. These neurons coincidently peaked at the reward generating time steps (.i.e., every fourth time steps).
Internalized time-keeping in the artificial agent
To ensure that the agent was not merely capturing visual characteristics of the training video, but had instead learned a general timing mechanism to produce the target interval, we tested it on different types of videos (see Methodology). This evaluation is analogous to determining whether humans are using specific visual cues in a video to mark the target interval, rather than estimating duration based on an internal sense of time.
Interestingly, the agent was able to successfully perform the time production task even when presented with a blank or constant frame video. Instead, its performance was slightly affected for video sequences with a higher MAFD (see Table 1). Thus, the agent successfully learns a general time-keeping mechanism and does not use visual characteristics of its input video as markers. Furthermore, it internalizes this time-keeping mechanism such that it performs optimally with minimal information from its environment.
The oscillating patterns in the LSTM hidden state activations which are supposed to have emerged as the agent’s time-keeping mechanism, were preserved across all test video sequences. The FFT analysis (not shown here) of the hidden state activations from all test video sequences indicated the same oscillation frequencies as in Figure 2 but with different magnitudes.
In order to check whether the agent relies on visual features of its input during training to learn the time production task, we trained two separate DRL agents (keeping the architecture and training iterations same as the original) with blank video frames and constant video frames, respectively. It is worth noting that these agents failed to learn the timing task. This indicates that the agent uses its visual input to extract temporal information during training. Thus, the time-keeping mechanism developed by the agent through interactions with its environment is eventually internalized and is minimally affected by changes in the environment.
A potentially related phenomena can be observed in the evolution of circadian rhythms in biological systems. Circadian clocks in organisms are believed to have developed as a strategy of photosynthesizing organisms to cope with day and night fluctuations from the environment Hut and Beersma, (2011). These organisms developed oscillating molecular behaviors (typically with periods close to 24 hours) in order to develop different molecular processes corresponding to daylight and darkness. Notably, in some species, these circadian rhythms persist even in the absence of external cues, such as in constant darkness Seki and Ito, (2022). This highlights how internal mechanisms, initially shaped by environmental interactions in biological systems, can become self-sustaining. We hypothesize that, the DRL agent’s behavior related to first learning a task by interacting with its environment, and later internalizing it to be able to perform the task irrespective of the environments input has some parallels with the evolution of biological processes like circadian clocks.
While the evolution of biological processes such as circadian rhythms involves highly complex mechanisms, the behavior exhibited by the DRL agent emerges from a much simpler setup. Therefore, these processes should not be considered directly analogous. Rather, the aim of this discussion is to propose hypotheses suggesting potential correlations between the behavior of a DRL agent performing a basic timing task and certain features of biological systems. Further evidence is needed to substantiate these correlations and to better understand their implications.
Average Frame Difference Average Reward
0.00 5.00
3.38 4.99
4.85 4.96
12.11 4.76
4.60 4.69
4.35 4.69
4.05 4.66
5.16 4.59
Table 1: Reward across videos with different levels of Mean Absolute Frame Difference (MAFD) (Salih and George,, 2020) averaged across all frames of the video sequence. The maximum achievable reward per episode is 5. While the agent generally performs well, its performance is slightly affected for video sequences with higher MAFD.
Discussion
The DRL agent learned to perform the duration production task and exhibited neural patterns and behaviors that are comparable to those observed in biological systems and biologically plausible models of timing. First, it converged to an internal state with oscillatory neural activations, a ubiquitous pattern in biological neurons. Second, its action network majorly relied on high amplitude oscillating neurons that coincidentally peaked at reward generating time steps, akin to the coincidence detection mechanism in the SBF model of timing. Third, the agent initially learned its time-keeping mechanism through embodied interactions but later internalized it, enabling it to perform the timing task with different environmental inputs. This process parallels the evolution of biological mechanisms such as circadian rhythms, where an organism builds an internal time representation through interactions with the environment and ultimately uses it as a buffer to adapt to environmental changes
The findings from this study, in conjunction with those of Deverett et al., (2019); Lin et al., (2023), suggest several implications for interval timing tasks. Specifically, the temporal representation required for successfully completing these tasks appears to be extractable by an artificial neural network from its input. This could imply the involvement of lower-level cognitive processes (König et al.,, 2013) in the execution of these tasks. This observation further supports the distinction made in time perception between duration judgment (i.e., the ability to estimate or reproduce time intervals) and the passage of time (i.e., how time feels) (Thönes and Stocker,, 2019). Specifically, while an artificial agent can learn to perform interval timing tasks with high accuracy, evaluating the subjective passage of time (whether time feels fast or slow), remains a significantly more challenging problem for such systems. This disparity may indicate the differing levels of cognitive complexity involved in the two tasks.
This study has several limitations. The parallels drawn between the agent’s behavior and biological systems, namely the resemblance to the SBF model of timing and the evolution of circadian rhythms, are preliminary and require further investigation. Given the complexity of biological systems, such analogies may overlook critical nuances. Two distinct systems arriving at similar solutions may not always imply shared mechanisms. This is because, the same problem can often be solved through multiple, distinct optimization pathways. Additionally, our analysis focused solely on the action network, whereas previous work has examined the value network and reported ramping-cell-like activity for temporal tasks (Lin et al.,, 2023). Future work could extend this study by analyzing the value network in our agent to uncover additional biologically plausible temporal patterns in the artificial neurons. Finally, this work represents an initial step towards using artificial agents to study temporal processing in biological systems. Understanding the implications of these similarities and translating them into meaningful biological insights remains a long-term challenge.
Acknowledgements
This work was supported by European Union’s Horizon 2020 FET research program under grant agreement No. 964464 (ChronoPilot).